# atomic orbital API Application Programmer Interface

### v1

@@ 7. Excited states
With the additional constrain, corresponding to the orthogonality of $\Psi$ to the ground state $\Psi_{\text{ground}}$, the new Lagrangian takes the form
@@
with symmetric matrix coefficients $\varepsilon_{ij}$. It is a function of $\varphi_{0},\ldots,\varphi_{M},\varepsilon,c_{0},\ldots,c_{M},$ $\varepsilon_{ij}$ with $i\leqslant j$ and $\lambda$. We calculate its gradient as follows.
$\frac{\delta\mathcal{L}}{\delta\varphi_{k}}$ $=c_{k}^{2}h\varphi_{k}+c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\varepsilon_{km}\varphi_{m}-\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}},\quad k=0,\ldots,M.$ (7.1)
$\frac{\partial\mathcal{L}}{\partial\varepsilon}$ $=-\frac{1}{4}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right).$ (7.2)
$\frac{\partial\mathcal{L}}{\partial c_{k}}$ $=\frac{1}{2}\sum_{m=0}^{M}H_{km}c_{m}-\frac{1}{2}\varepsilon c_{k}-\frac{1}{2}\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2},\quad k=0,\ldots,M.$ (7.3)
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}$ $=-\frac{1}{2}\left(\left\|\varphi_{k}\right\|^{2}-1\right),\quad k=0,\ldots,M,$ (7.4)
$\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}$ $=-\int\varphi_{i}\varphi_{j},\quad 0\leqslant i<j\leqslant M.$ (7.5)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=-\frac{1}{2}\sum_{m,n=0}^{M,M^{\text{gr}}}c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)^{2}$ (7.6)
Now let us calculate the Hessian $d\nabla\mathcal{L}$.
$\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\varphi_{j}=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ -\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}},$ (7.7)
where $k=0,\ldots,M.$
$\frac{\partial}{\partial\varepsilon}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}=0$
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta c_{j}=2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\lambda\sum_{n=0}^{M^{\text{gr}}}\delta c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}$
@@
where we extended the orbital energy update by symmetry.
$\frac{\partial}{\partial\lambda}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\lambda=-\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}$
Summing these equalities we obtain
$d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ +2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m}\\ -\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-\lambda\sum_{n=0}^{M^{\text{gr}}}\delta c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}-\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}$ (7.8)
$d\frac{\partial\mathcal{L}}{\partial\varepsilon}\delta w=-\frac{1}{2}\sum_{m=0}^{M}c_{m}\delta c_{m}$ (LHS2)
@@
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta c_{j}=\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}$
$\frac{\partial}{\partial{\lambda}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta\lambda=-\frac{1}{2}\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2}$
Summing these equalities we obtain
$d\frac{\partial\mathcal{L}}{\partial c_{k}}\delta w=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km)-\frac{1}{2}\delta\varepsilon c_{k}+\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k}\\ -\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-\frac{1}{2}\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2}$ (7.9)
Finally,
$d\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}\delta w=-\int\varphi_{k}\delta\varphi_{k},\quad k=0,\ldots,M,$ (LHS4)
$d\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}\delta w=-\int\delta\varphi_{i}\varphi_{j}-\int\varphi_{i}\delta\varphi_{j},\quad 0\leqslant i<j\leqslant M,$ (LHS5)
$d\frac{\partial\mathcal{L}}{\partial\lambda}\delta w=-\sum_{m,n=0}^{M,M^{\text{gr}}}c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{m}\varphi_{n}^{\text{gr}}-\frac{1}{2}\sum_{m,n=0}^{M,M^{\text{gr}}}\delta c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)^{2}.$ (LHS6)
The right hand side is $-\nabla\mathcal{L}$. This finishes the detailed description of the Newton equation. Now we want to rewrite it in the self-consistent form, which will be suitable for an iterative procedure.
@@ 7.1. Self-consistent form
is fixed at each Newton step, as follows.
@@
$\varepsilon-H$ is an $(M+1)\times(M+1)$ matrix. Here
$f_{k}=\sum_{m=0}^{M}H_{km}c_{m}-\varepsilon c_{k}-\lambda v_{k}+4c_{k}(\delta k|h|k)+2\sum_{m=0}^{M}c_{m}(\delta km|km)+2\sum_{m=0}^{M}c_{m}(k\delta m|km)\\ -2\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{k}\varphi_{n}^{\text{gr}},\quad k=0,\ldots,M$
@@
as follows
$\left(-\frac{c_{k}^{2}}{2}\Delta-\varepsilon_{kk}\right)\delta\varphi_{k}+\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\left(-\frac{c_{k}^{2}}{2}\Delta-\varepsilon_{kk}\right)\varphi_{k}+\left(\frac{2\delta c_{k}}{c_{k}}\varepsilon_{kk}-\delta\varepsilon_{kk}\right)\varphi_{k}\\ +c_{k}^{2}V_{\text{nuc}}\left(\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\varphi_{k}+\delta\varphi_{k}\right)-\sum_{\begin{subarray}{c}m=0\\ m\neq k\end{subarray}}^{M}\left(\varepsilon_{km}\delta\varphi_{m}+\delta\varepsilon_{km}\varphi_{m}+\varepsilon_{km}\varphi_{m}\right)\\ +\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}J(km)(\varphi_{m}+\delta\varphi_{m})\\ +c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}\right)\\ -\lambda c_{k}\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-(\lambda\delta c_{k}+\delta\lambda c_{k}+\lambda c_{k})\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}=0$
Finally, we arrive to (6.8) with
@@
and the convolution operator $R_{k}$ is defined as above. This finishes the description of $\mathbb{F}_{k}$ in (6.7).
@@ 7.2. Numerics
Figure 6. Three determinants excited state of the hydrogen molecule with the coefficients $c_{0}=0.76103,c_{1}=-0.64796$ and $c_{2}=-0.03128$, accordingly. The corresponding energy is $-0.68989$.
For He atom high accuracy excited states were computed in [3]. The first excited singlet state has the energy $-2.14597404605441741580502897546192$, that serves as a reference. Impressively, our method gives the energy $-2.14350$ for two determinants associated to coefficients $c_{0}=0.75880,c_{1}=-0.65132$ and $M^{\text{gr}}=5$ corresponding to 6 configurations in the ground state.
For H2 molecule we calculate the first closed shell excited state at the same internuclear distance $R_{\text{nuc}}=1.4010784$ as above. Using 6 configurations in the ground state for the constrain, we report the energy $-0.68844$ for two configurations with $c_{0}=0.76135,c_{1}=-0.64834$ and the energy $-0.69028$ for 5 determinants with $c_{0}=0.76125,c_{1}=-0.64732,c_{2}=-0.03140,c_{3}=0.01669,c_{4}=0.01461$. The orbitals can be observed in Figure 6. In this case we do not have a high precision energy value, which one could use as a reference, but some close values were reported in [7, 18].

### v2

@@ 6. Excited states
With the additional constraint, corresponding to the orthogonality of $\Psi$ to the ground state $\Psi_{\text{ground}}$, the new Lagrangian takes the form
@@
with symmetric matrix coefficients $\varepsilon_{ij}$. It is a function of $\varphi_{0},\ldots,\varphi_{M},\varepsilon,c_{0},\ldots,c_{M},$ $\varepsilon_{ij}$ with $i\leqslant j$ and $\lambda$. Its gradient $\nabla\mathcal{L}$ consists of the following components
(6.1) $\frac{\delta\mathcal{L}}{\delta\varphi_{k}}$ $=c_{k}^{2}h\varphi_{k}+c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\varepsilon_{km}\varphi_{m}-\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}},\quad k=0,\ldots,M,$
(6.2) $\frac{\partial\mathcal{L}}{\partial\varepsilon}$ $=-\frac{1}{4}\left(\sum_{m=0}^{M}c_{m}^{2}-1\right),$
(6.3) $\frac{\partial\mathcal{L}}{\partial c_{k}}$ $=\frac{1}{2}\sum_{m=0}^{M}H_{km}c_{m}-\frac{1}{2}\varepsilon c_{k}-\frac{1}{2}\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2},\quad k=0,\ldots,M,$
(6.4) $\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}$ $=-\frac{1}{2}\left(\left\|\varphi_{k}\right\|^{2}-1\right),\quad k=0,\ldots,M,$
(6.5) $\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}$ $=-\int\varphi_{i}\varphi_{j},\quad 0\leqslant i<j\leqslant M,$
(6.6) $\frac{\partial\mathcal{L}}{\partial\lambda}$ $=-\frac{1}{2}\sum_{m,n=0}^{M,M^{\text{gr}}}c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)^{2}.$
Next we compute the Hessian $d\nabla\mathcal{L}$. We start by differentiating the variational derivatives. First, with respect to orbitals
(6.7) $\sum_{j=0}^{M}\partial_{\varphi_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\varphi_{j}=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ -\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}},$
where $k=0,\ldots,M.$ Then with respect to Lagrange multiplier $\varepsilon$ as
$\frac{\partial}{\partial\varepsilon}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}=0.$
The derivatives with respect to CI coefficients equal
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta c_{j}=2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\lambda\sum_{n=0}^{M^{\text{gr}}}\delta c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}.$
The derivatives with respect to orbital energies are
@@
where we extended the orbital energy update by symmetry. Finally, the derivative with respect to Lagrange multiplier $\lambda$ responsible for the orthogonality to the ground state has the following expression
$\frac{\partial}{\partial\lambda}\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta\lambda=-\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}.$
Summing these equalities we obtain the first component
(6.8) $d\frac{\delta\mathcal{L}}{\delta\varphi_{k}}\delta w=c_{k}^{2}h\delta\varphi_{k}-\sum_{m=0}^{M}\varepsilon_{km}\delta\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}+J(km)\delta\varphi_{m}\right)\\ +2c_{k}\delta c_{k}h\varphi_{k}+\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}-\sum_{m=0}^{M}\delta\varepsilon_{km}\varphi_{m}\\ -\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-\lambda\sum_{n=0}^{M^{\text{gr}}}\delta c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}-\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{k}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}$
of the Hessian $d\nabla\mathcal{L}(w)$ at the update $\delta w$. The second component is given by
(LHS2) $d\frac{\partial\mathcal{L}}{\partial\varepsilon}\delta w=-\frac{1}{2}\sum_{m=0}^{M}c_{m}\delta c_{m}.$
In order to obtain the third component, we compute the variations
@@
with respect to orbitals and the partial derivatives
$\frac{\partial}{\partial{\varepsilon}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta\varepsilon=-\frac{1}{2}\delta\varepsilon c_{k},$
$\sum_{j=0}^{M}\frac{\partial}{\partial c_{j}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta c_{j}=\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}-\frac{1}{2}\varepsilon\delta c_{k},$
$\frac{\partial}{\partial{\lambda}}\frac{\partial\mathcal{L}}{\partial c_{k}}\delta\lambda=-\frac{1}{2}\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2}.$
Summing these equalities we obtain the third component
(6.9) $d\frac{\partial\mathcal{L}}{\partial c_{k}}\delta w=2c_{k}(\delta k|h|k)+\sum_{m=0}^{M}c_{m}(\delta km|km)+\sum_{m=0}^{M}c_{m}(k\delta m|km)-\frac{1}{2}\delta\varepsilon c_{k}+\frac{1}{2}\sum_{m=0}^{M}H_{km}\delta c_{m}\\ -\frac{1}{2}\varepsilon\delta c_{k}-\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-\frac{1}{2}\delta\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)^{2}.$
Finally, we compute the remaining components
(LHS4) $d\frac{\partial\mathcal{L}}{\partial\varepsilon_{kk}}\delta w=-\int\varphi_{k}\delta\varphi_{k},\quad k=0,\ldots,M,$
(LHS5) $d\frac{\partial\mathcal{L}}{\partial\varepsilon_{ij}}\delta w=-\int\delta\varphi_{i}\varphi_{j}-\int\varphi_{i}\delta\varphi_{j},\quad 0\leqslant i<j\leqslant M,$
(LHS6) $d\frac{\partial\mathcal{L}}{\partial\lambda}\delta w=-\sum_{m,n=0}^{M,M^{\text{gr}}}c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{m}\varphi_{n}^{\text{gr}}-\frac{1}{2}\sum_{m,n=0}^{M,M^{\text{gr}}}\delta c_{m}c_{n}^{\text{gr}}\left(\int\varphi_{m}\varphi_{n}^{\text{gr}}\right)^{2}.$
Expressions (6.8)-(LHS6) define $d\nabla\mathcal{L}(w)(\delta w)$ staying at the left-hand side of the Newton equation. The right hand side is $\nabla\mathcal{L}(w)$ with the gradient components (6.1)-(6.6). This finishes the detailed description of the Newton equation. Next we rewrite it in the self-consistent form, which will be suitable for an iterative procedure.
@@ 6.1. Self-consistent form
is fixed at each Newton step, as follows. First, we separate evaluation of $\delta\varepsilon,\delta\lambda,\delta\mathbf{c}$ as
@@
$\varepsilon-H$ is an $(M+1)\times(M+1)$ matrix. The column vector $f$ has the components
$f_{k}=\sum_{m=0}^{M}H_{km}c_{m}-\varepsilon c_{k}-\lambda v_{k}+4c_{k}(\delta k|h|k)+2\sum_{m=0}^{M}c_{m}(\delta km|km)+2\sum_{m=0}^{M}c_{m}(k\delta m|km)\\ -2\lambda\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\left(\int\varphi_{k}\varphi_{n}^{\text{gr}}\right)\int\delta\varphi_{k}\varphi_{n}^{\text{gr}},\quad k=0,\ldots,M.$
@@
by inverting the kinetic energy in the equations
$\left(-\frac{c_{k}^{2}}{2}\Delta-\varepsilon_{kk}\right)\delta\varphi_{k}+\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\left(-\frac{c_{k}^{2}}{2}\Delta-\varepsilon_{kk}\right)\varphi_{k}+\left(\frac{2\delta c_{k}}{c_{k}}\varepsilon_{kk}-\delta\varepsilon_{kk}\right)\varphi_{k}\\ +c_{k}^{2}V_{\text{nuc}}\left(\left(1+\frac{2\delta c_{k}}{c_{k}}\right)\varphi_{k}+\delta\varphi_{k}\right)-\sum_{\begin{subarray}{c}m=0\\ m\neq k\end{subarray}}^{M}\left(\varepsilon_{km}\delta\varphi_{m}+\delta\varepsilon_{km}\varphi_{m}+\varepsilon_{km}\varphi_{m}\right)\\ +\delta c_{k}\sum_{m=0}^{M}c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}\delta c_{m}J(km)\varphi_{m}+c_{k}\sum_{m=0}^{M}c_{m}J(km)(\varphi_{m}+\delta\varphi_{m})\\ +c_{k}\sum_{m=0}^{M}c_{m}\left(J(\delta km)\varphi_{m}+J(k\delta m)\varphi_{m}\right)\\ -\lambda c_{k}\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\delta\varphi_{k}\varphi_{n}^{\text{gr}}-(\lambda\delta c_{k}+\delta\lambda c_{k}+\lambda c_{k})\sum_{n=0}^{M^{\text{gr}}}c_{n}^{\text{gr}}\varphi_{n}^{\text{gr}}\int\varphi_{k}\varphi_{n}^{\text{gr}}=0.$
Finally, we arrive to (5.9) with the new
@@
and the convolution operator $R_{k}$ is defined as above. This finishes the description of $\mathbb{F}_{k}$ in (5.8) for the first excited state problem.
@@ 6.2. Numerics
Figure 6. Three determinants excited state of the hydrogen molecule with the coefficients $c_{0}=0.76103,c_{1}=-0.64796$ and $c_{2}=-0.03128$, accordingly. The corresponding energy is 0.68989-0.68989.
For He atom high accuracy excited states were computed in [1]. The first excited singlet state has the energy 2.14597404605441741580502897546192-2.14597404605441741580502897546192, that serves as a reference. Impressively, our method gives the energy 2.14350-2.14350 for two determinants associated to coefficients $c_{0}=0.75880,c_{1}=-0.65132$ and $M^{\text{gr}}=5$ corresponding to 6 configurations in the ground state.
For H2 molecule we calculate the first closed shell excited state at the same internuclear distance $R_{\text{nuc}}=1.4010784$ as above. Using 6 configurations in the ground state for the constrain, we report the energy 0.68844-0.68844 for two configurations with $c_{0}=0.76135,c_{1}=-0.64834$ and the energy 0.69028-0.69028 for 5 determinants with $c_{0}=0.76125,c_{1}=-0.64732,c_{2}=-0.03140,c_{3}=0.01669,c_{4}=0.01461$. The orbitals can be observed in Figure 6. In this case we do not have a high precision energy value, which one could use as a reference, but some close values were reported in [4, 19]. Figure 6 cannot convey the orthogonality of the orbitals $\varphi_{0}$ and $\varphi_{1}$, since, to a large extend, their overlap is canceled by contributions from the long tails.